OpenStudy (anonymous):

Solve 6 and 2/9 - 4 and 1/3. Show your work

OpenStudy (anonymous):

6 and 2/9 - 4 and 1/3 the first thing I would do is change them into mixed numbers 56/9 - 13/3 Find a common denominator (9 in this case) 56/9 - 39/9 = 17/9 = 1 and 8/9

OpenStudy (anonymous):

do you mean \[6+2/9-4+1/3\]?

OpenStudy (anonymous):

yes